OpenStudy (anonymous):

You deposit $750 in an account that earns 7.5% simple interest. How much will be in the account after 5 years?

OpenStudy (anonymous):

A=P(1+r)^t A=750(1+0.075)^5 A=1067.721995≈1,067.72 I think so, you must check! I'm not sure

OpenStudy (anonymous):

@dalia_lam It's called simple interest. For an example according to my paper, it says: You deposit $750 in an account that earns 7.5% simple interest. How much will be in the account after 5 years? Answer: 200 x .05 (5%) - 10 20 x 10 = 200 200 current dollars + interest dollars = 400 dollars.

OpenStudy (anonymous):

sorry, i thought it was compound interest

OpenStudy (anonymous):

I'm sorry, I meant: *200 x .05 (5%) = 10

OpenStudy (anonymous):

I just want to say that I like your picture!